• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Moffit

    Log entry time 08:06:08 on September 18, 2005

    Entry number 155079

    keyword=Owl Shift Summary

    Owl Shift for Sunday September 18th, 2005


    Leader:   Bryan Moffit
    Worker/Target:   Brandon Craver

    Left HRS p0   = -2.737 GeV/c   Septum Set Current = 283.57A
    Right HRS p0  = -2.733 GeV/c   Septum Set Current = 278.22A
    Sweeper is ON at 111.00A

    Summary

    Mostly production (production runs 5136-5147: 284.0k pairs ~= 0.76C). Spent a couple of hours doing some profile scans at different spectrometer momentum settings.

    Significant Events

    • 00:00 Began shift with the start of run 5136, compton run 12207. IHWP IN.
    • 00:20 Asked MCC to increase beam current to 40µA (from 38µA).
    • 03:44 Begin profile scans at different momentum settings.
    • 03:56 Cycle Q2 and Q3 to prepare for higher momentum runs.
    • 04:32 Noticed that the right arm Q1 would not change to the new momentum settings. Ignored it.
    • 05:00 Right Q1 dropped its current to zero (slowly). Reset it with the posted procedure, and set its current by hand.
    • 05:38 Return magnets to nominal setpoints.
    • 05:40 Incoming Call: R. Suleiman: Status Update
    • 05:51 Asked MCC to retract the IHWP. Back to production.
    • 07:30 Compton DAQ crashes hard. ROCs will not respond to reboot.

    • Data Run Summary

      • 00:00 Run 5136 - Production Run (20cm Helium-4), IHWP IN, 48.9k pairs.
      • 00:55 Run 5137 - Production Run (20cm Helium-4), IHWP IN, 46.7k pairs.
      • 01:51 Run 5138 - Production Run (20cm Helium-4), IHWP IN, 47.9k pairs.
      • 02:47 Run 5139 - Production Run (20cm Helium-4), IHWP IN, 48.0k pairs.
      • 03:44 Run 5140 - Profilescan (20cm Helium-4), IHWP IN. 0% of nominal momentum.
      • 04:14 Run 5141 - Profilescan (20cm Helium-4), IHWP IN. +2% of nominal momentum. Right arm still ramping.
      • 04:37 Run 5142 - Profilescan (20cm Helium-4), IHWP IN. +1% of nominal momentum. Right Q1 wrong.
      • 04:52 Run 5143 - Profilescan (20cm Helium-4), IHWP IN. +0.5% of nominal momentum. Right Q1 wrong.
      • 05:14 Run 5144 - Profilescan (20cm Helium-4), IHWP IN. -10MeV of nominal momentum. Right Q1 wrong.
      • 05:39 Run 5145 - Profilescan (20cm Helium-4), IHWP IN. 0% of nominal momentum.
      • 05:54 Run 5146 - Production Run (20cm Helium-4), IHWP OUT, 47.8k pairs.
      • 06:50 Run 5147 - Production Run (20cm Helium-4), IHWP OUT, 44.7k pairs.
      • 07:46 Run 5148 - Production Run (20cm Helium-4), IHWP OUT. Passed to day shift.

      Compton Data Run Summary

      • 00:00 Run 12208 - IHWP IN. ROC2 in disconnected state. Barely any coincidence rate with cavity on.
      • 00:35 Run 12209 - IHWP IN, Beam polarization = -82.5%,
      • 02:39 Run 12210 - IHWP IN, Beam polarization = -82.4%,
      • 05:46 Run 12211 - IHWP OUT. ROC5 crashed. CODA rebooted, but ROCS will not reboot.